OpenStudy (anonymous):

hi you guys, I need serious help with problems.:) What is the range of the function f(x) = -10x for the domain {-4, -2, 0, 2, 4}? a.{40, 20, 0, -20, -30} b.{0, 20, 40, 60, 80} c.{-4, -2, 0, 2, 4} d.{-40, -20, 0, 20, 40}

OpenStudy (anonymous):

d. Put the values of x=-4,-2,0,2,4 in the eqn. f(x)=-10*x, and you get the set of values.

OpenStudy (anonymous):

thanks so much .. can you help with a few more

OpenStudy (anonymous):

The number of hours required to do a painting job is represented by the function f(x)=2x , where f(x) gives the number of hours and x is the number of painters doing the job. How much time will {4, 10, 14, 16, 18} workers take to do the job? a.{0.2, 0.7, 0.8, 0.9, 1.0} b.{0.1, 0.4, 0.6, 0.8, 0.9} c.{0.2, 0.59, 0.73, 0.84, 0.9} d.{0.11, 0.125, 0.14, 0.2, 0.5}

OpenStudy (anonymous):

do i divide or multiply when i try to multiply i do not get a decimal

OpenStudy (anonymous):

I don't get this question...

OpenStudy (anonymous):

i have to plug the range of numbers in f(x) i think because it says f(x) gives number of hours and the question is how many hours will it take

OpenStudy (perl):

"The number of hours required to do a painting job is represented by the function f(x)=2x , where f(x) gives the number of hours and x is the number of painters doing the job" This function does not make sense. the more painters you have, the less time it will take to finish the painting job.

OpenStudy (perl):

f(x) = 2 / x makes more sense

OpenStudy (anonymous):

so i am dividing the range by 2?

OpenStudy (anonymous):

i finished the answer already thanks :) idk how u guys do this

OpenStudy (anonymous):

If the range of the function f(x)= 5x − 3 is {-13, -8, -3, 12, 17}, what is its domain?\ {0, 1, 2, 3, 4} {-2, -1, 0, 1, 2} {-2, -1, 0, 3, 4} {-4, -3, 0, 3, 4}

OpenStudy (anonymous):

do i plug in -13 for (x)

OpenStudy (anonymous):

and so on....

OpenStudy (anonymous):

@perl :Exactly what I was thinking. 2/x could be possible. Otherwise, this is a wrong function, no question.